# HiFICL: High-Fidelity In-Context Learning for Multimodal Tasks

### 13 Mar 2026 v1

@@ Abstract
In-Context Learning (ICL) is a significant paradigm for Large Multimodal Models (LMMs), using a few in-context demonstrations (ICDs) for new task adaptation. However, its performance is sensitive to demonstration configurations and computationally expensive. Mathematically, the influence of these demonstrations can be decomposed into a dynamic mixture of the standard attention output and the context values. Current approximation methods simplify this process by learning a ”shift vector”. Inspired by the exact decomposition, we introduce High-Fidelity In-Context Learning (HiFICL) to more faithfully model the ICL mechanism. HiFICL consists of three key components: 1) a set of ”virtual key-value pairs” to act as a learnable context, 2) a low-rank factorization for stable and regularized training, and 3) a simple end-to-end training objective. From another perspective, this mechanism constitutes a form of context-aware Parameter-Efficient Fine-Tuning (PEFT). Extensive experiments show that HiFICL consistently outperforms existing approximation methods on several multimodal benchmarks. The code is available at https://github.com/bbbandari/HiFICL.
Accepted to the IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR) 2026
@@ 1 Introduction
To mitigate these issues, a dominant research direction has focused on approximating the ICL effect. This paradigm, centered on learning a “shift vector,” aims to distill the knowledge from explicit ICDs into a compact representation that can be efficiently injected into the model [14, 38, 34, 17, 24]. While this approach has proven effective in reducing inference costs, we argue that the entire paradigm is predicated on a theoretically imprecise and ultimately limiting assumption. These methods treat the “shift” as an external, additive component to be learned, focusing on approximating the emergent outcome of the ICL process. This overlooks a more fundamental question: what is the underlying causal mechanism that generates this effect in the first place?
@@
Figure 1: Overview of the HiFICL framework. (a) The overall training process involves a frozen LMM backbone and layer-wise trainable virtual key-value pairs, optimized via a final task loss. (b) Inside each attention head, these virtual pairs are used to compute the high-fidelity attention output according to our derived formula (Eq. 1). (c) Conceptually, our dynamic, query-dependent approximation (yellow line) provides a more faithful fit to the underlying data manifold compared to previous linear shift approximations (blue line).
This reframing of the ICL approximation problem aligns with a growing body of research questioning the adequacy of the linear shift model. Recent studies have begun to theorize that ICL itself functions as a form of ”inference-time finetuning,” where the model’s behavior is dynamically and temporarily optimized by the provided context, much like a few steps of gradient descent [7, 21]. This perspective, supported by empirical evidence showing a striking similarity between the effects of ICL and finetuning on model confidence [2], suggests that the ICL mechanism is fundamentally a learning process, not a simple vector addition. While these works provide the theoretical and empirical grounding for this hypothesis, our work provides the first concrete instantiation of it as a training-time PEFT method. HiFICL is designed to distill and solidify this transient, inference-time adaptation into a permanent and efficient set of trainable parameters. Viewed through this lens, its mechanism presents a more logically coherent form of adaptation than conventional methods like LoRA [15], which apply static, input-agnostic updates. By directly simulating the foundational process of ICL through learnable, context-aware modulation, HiFICL constitutes a more powerful and principled form of adaptation.
@@ 2.1 ICL in Large Multimodal Models
In-Context Learning (ICL) has become a key capability of Large Language Models (LLMs) [4, 9], enabling few-shot task adaptation without updating model parameters. This powerful paradigm has been successfully extended to the multimodal domain in Large Vision-Language Models (LVLMs). Pioneering models like Flamingo [1], OpenFlamingo [3], and IDEFICS [18] have demonstrated remarkable few-shot abilities by conditioning on interleaved image-text sequences, showcasing the potential of multimodal ICL [39].
However, integrating ICL into the multimodal space introduces unique and exacerbated challenges compared to the text-only domain [31]. First, the high token cost of visual inputs drastically limits the context window and increases computational overhead, making the inclusion of multiple high-resolution image demonstrations computationally prohibitive [28, 44]. Second, the synergistic effects between vision and language modalities amplify the sensitivity of ICL performance to the selection and ordering of demonstrations [23, 30]. Strategies that are effective in LLMs, such as similarity-based retrieval, can sometimes perform worse than random selection in LMMs, as visual similarity may not always align with task-relevant semantics [46]. This highlights the outstanding challenge of identifying optimal demonstration configurations in the complex multimodal feature space [35, 47]. While some works have focused on mitigating this instability by optimizing the demonstration sequence [26] or modulating the internal attention mechanism [25, 22], the fundamental hurdles of high computational cost and extreme sensitivity remain. These challenges together strongly motivate the paradigm of ICL approximation, which our work aims to advance.
@@ 2.2 The Linear Shift Approximation
However, recent advances in mechanistic interpretability reveal a fundamental contradiction with this linear paradigm. Research has shown that ICL is not a simple global shift but is executed by specialized circuits. A key discovery is the role of “Induction Heads” [33, 11], attention heads that perform complex pattern matching and retrieval. These mechanisms are believed to be how models implicitly implement learning algorithms like gradient descent [7, 43]. Furthermore, analyses from the perspective of hidden state geometry demonstrate that ICL is a highly non-linear transformation that dynamically reshapes the representation space to enhance class separability and align with decision boundaries [45]. This stark discrepancy between the linear assumption of current approximation methods and the non-linear reality of the underlying mechanism suggests that the linear shift paradigm itself is a theoretical bottleneck, motivating the exploration of more faithful approximation models.
@@ 2.3 Low-Rank Adaptation for PEFT
The connection between In-Context Learning (ICL) and finetuning is not merely a conceptual analogy but is rooted in the mechanistic underpinnings of the Transformer architecture. A growing body of theoretical research posits that the self-attention mechanism can implicitly simulate optimization algorithms. These studies show that attention heads can identify relevant context examples and aggregate their label information in a manner that mathematically mirrors one or more steps of gradient descent on a prediction loss [7, 21, 36]. This frames ICL as a form of ”inference-time” optimization, where the model effectively fine-tunes its internal representations dynamically for each query.
This perspective naturally raises a question: how does this implicit, dynamic optimization relate to explicit, parameter-based finetuning? Parameter-Efficient Fine-Tuning (PEFT) [13], and its most prominent method, Low-Rank Adaptation (LoRA) [15], represent the dominant paradigm for the latter. LoRA injects trainable, low-rank matrices into the frozen layers of a pre-trained model, offering a computationally efficient alternative to full finetuning [8]. However, a core characteristic of LoRA is that its adaptation is static and input-agnostic. This creates a conceptual gap: while ICL performs a dynamic, per-query optimization, LoRA performs a static, global one. Our work is motivated by bridging this gap, aiming to create a PEFT method that is not only parameter-efficient but also retains the dynamic, context-aware adaptation properties inherent to the ICL mechanism.
@@ 3.1 Mathematical Analysis
where $\text{SA}(q,K,V)$ is the standard self-attention output over the query tokens and $V_{D}$ can be viewed as the redefined offset vector basis. The coefficients $\alpha(q)$ and $\beta(q)$ are dynamically computed scalar and vector weights, respectively, defined as:
$\alpha(q)=\frac{Z_{2}}{Z_{1}+Z_{2}},\quad\beta(q)=\frac{1}{Z_{1}+Z_{2}}\exp\left(\frac{qK_{D}^{\top}}{\sqrt{d_{k}}}\right)$ (2)
@@ 3.2 HiFICL: Direct Parameterization
Based on the theoretical insights from Sec. 3.1, we introduce HiFICL. The core idea is to directly parameterize the source of ICL, the unknown demonstration key-value pairs $(K_{D},V_{D})$, rather than approximating its ultimate effect. We achieve this by introducing a set of learnable soft prompts, which manifest as “virtual key-value pairs,” $(K_{learn},V_{learn})$, into the attention mechanism. To capture the functional specialization of different attention heads [6, 42], our modifications are applied on a per-head basis. We equip each attention head $h$ with an independent set of $n$ virtual pairs, $(K_{learn}^{(h)},V_{learn}^{(h)})$, allowing each head to learn the specialized contextual information it requires for its sub-task.
@@
where $n$ is the number of learnable prompts, $K_{A}^{(h)},V_{A}^{(h)}\in\mathbb{R}^{n\times r}$ and $K_{B}^{(h)},V_{B}^{(h)}\in\mathbb{R}^{r\times d_{h}}$, with the rank $r\ll d_{h}$. This design elegantly provides two critical benefits. First, for structural stability, we initialize the $V_{B}^{(h)}$ matrices to zero. This guarantees that the contextual shift term in Eq. (1) is null at the beginning of training, preventing potential gradient explosion and creating a smooth learning trajectory from the base model’s state. Second, the low-rank factorization of $K_{learn}^{(h)}$ acts as a powerful form of structural regularization. It compels the model to learn a compact and generalizable set of “prototype keys,” effectively creating an information bottleneck that mitigates overfitting.
To maximize the potential of this high-fidelity architecture, we adopt a simple yet powerful teacher-free, end-to-end optimization strategy. We discard the complex and potentially restrictive teacher-student paradigm used in prior work [17], which relies on intermediate alignment losses. Instead, we optimize all trainable parameters, including the injected low-rank virtual matrices, using only the final supervised task loss. For generative tasks, this is typically the cross-entropy loss over the target sequence:
@@ 3.4 HiFICL as a Dynamic PEFT
Beyond ICL approximation, HiFICL introduces a new perspective on Parameter-Efficient Fine-Tuning (PEFT). Recent studies have begun to theorize that ICL itself functions as a form of ”inference-time finetuning,” where the model’s behavior is dynamically and temporarily optimized by the provided context [21, 2]. Our work provides the first concrete instantiation of this hypothesis as a training-time PEFT method. We situate HiFICL by contrasting it with Low-Rank Adaptation (LoRA) [15], a dominant technique that performs a static, input-agnostic adaptation in the weight space (Fig. 3). In stark contrast, HiFICL operates in the activation space, performing a dynamic, content-aware adaptation by providing the model with a learnable memory of virtual contextual examples. Its mechanism is thus more interpretable, analogous to teaching a model how to reason with context rather than applying a ”black-box” modification to its weights. As our experiments will demonstrate, this principled design, which effectively distills the ”inference-time finetuning” mechanism into a permanent set of parameters, not only leads to superior performance but also achieves significantly higher parameter efficiency.
@@ 4.1 Setup
To ensure a fair comparison, our implementation details for all trainable baselines strictly follow those of MimIC [17]. We employ the AdamW optimizer [29] with a learning rate of 5e-3, coupled with a cosine annealing scheduler with a warmup phase over the first 10% of training steps. For HiFICL, we set the number of learnable prompts to $n=8$ across all experiments. The rank $r$ is a task-dependent hyperparameter, with detailed configurations provided in the Appendix.
@@ 4.2 Comparison with Existing Methods
Compared methods.
We compare HiFICL against several categories of strong baselines. (1) In-Context Learning (ICL): We evaluate the performance of Zero-shot and 8-shot ICL with randomly selected demonstrations. (2) PEFT: We include LoRA [15], a widely-used PEFT method that fine-tunes the model by adding low-rank adapters to the attention weights. We apply LoRA with a rank of 16 to the attention matrices ($W_{q},W_{k},W_{v},W_{o}$) of the language model layers. (3) ICL Approximation: We benchmark against state-of-the-art methods from this paradigm which modify the language model part of the LMM. These methods differ in their insertion points: LIVE [34] introduces learnable vectors after each FFN layer, whereas both the prior state-of-the-art, MimIC [17], and our HiFICL operate directly within the Multi-Head Attention (MHA) block of each layer. For all trainable methods, we ensure a fair comparison by using the same training data and core hyperparameter settings.
Results analysis.
Tab. 1 presents the results of HiFICL compared to all baselines. For the ICL baselines, we observe that 8-shot ICL significantly outperforms Zero-shot, establishing a strong performance baseline. For the trainable methods, nearly all approaches consistently improve over the 8-shot ICL baseline, with HiFICL demonstrating a commanding advantage. On the LLaVA-Interleave-7b model, HiFICL surpasses all competing methods, establishing new state-of-the-art results across all tasks. Notably, it outperforms the prior best method, MimIC, by a substantial 1.9% on OK-VQA and 1.1% on COCO Captioning (CIDEr). This superiority is further amplified on the Idefics2-8b-base model, where the performance gap over MimIC widens to a remarkable 2.79% on VQAv2. It is particularly noteworthy that HiFICL achieves this with a modest parameter budget (1.1-4.4M). For instance, on LLaVA, it requires approximately 8x fewer parameters than LoRA and MimIC while delivering superior results. These robust empirical gains underscore the fundamental advantages of our high-fidelity, end-to-end design.
@@
We further analyze the learning efficiency of HiFICL by evaluating performance on five different data scales: 100, 200, 300, 500, and 1000 samples. As illustrated in Fig. 4, HiFICL consistently demonstrates strong data efficiency across both the LLaVA and Idefics2 models on VQAv2 and COCO benchmarks. In most scenarios, HiFICL exhibits a steeper learning curve than competing methods, particularly in the low-data regime (100-300 samples). For instance, on Idefics2 for COCO, HiFICL surpasses the strong 8-shot ICL baseline with only around 300 training samples. This superior sample efficiency suggests that HiFICL provides a more direct and effective learning signal, enabling the model to capture essential task semantics more rapidly.
@@ 4.3 Ablations and More Analyses
Core component ablations.
We systematically deconstruct HiFICL on the Idefics2 model to validate its key design principles, with results detailed in Tab. 2. The findings confirm the necessity of each component. First, re-implementing HiFICL within a teacher-student framework (+ Teacher) by aligning hidden states similar to MimIC [17] leads to a consistent performance drop (e.g., VQAv2 accuracy falls from 72.08% to 70.09%), supporting our hypothesis that the teacher model acts as a performance ceiling. Second, the dual low-rank decomposition is vital; removing the factorization from either the key matrix (- LoRA on K) or the value matrix (- LoRA on V) significantly degrades performance, confirming their respective roles as a structural regularizer and an essential stabilizer. Finally, disabling the self-attention scaling factor ($\alpha=1$), which reverts our method to a simpler linear shift approximation, also causes a substantial performance decline across all benchmarks. This empirically validates that capturing the full, non-linear attention dynamic is a key driver of HiFICL’s superiority.
@@
A primary motivation for ICL approximation is to reduce the high inference cost of long contexts. We evaluate this by comparing the average processing time per sample on VQAv2 using the LLaVA-Interleave-7b model. As shown in Fig. 6, all ICL approximation methods, including HiFICL, offer a dramatic speedup over standard multi-shot ICL. HiFICL’s inference runtime is nearly on par with the Zero-shot baseline and comparable to prior approximation methods. Crucially, it is approximately 1.8x faster than 8-shot ICL and 3.1x faster than 16-shot ICL, confirming that our high-fidelity approach retains the core efficiency benefits of ICL approximation.
Hyperparameter analysis: Task-adaptive rank.
To explore the properties of our low-rank decomposition, we analyze how the optimal rank $r$ varies with task complexity. We evaluate HiFICL on Idefics2 using different ranks for both VQAv2 and the more challenging, knowledge-intensive OK-VQA. As illustrated in Fig. 7, the two datasets exhibit distinct optimal rank configurations. For VQAv2, performance peaks at a rank of $r=8$ and then declines, suggesting that a higher rank may lead to overfitting on this relatively simpler task. In contrast, for the more complex OK-VQA, performance continues to improve up to a higher rank of $r=16$ before dropping. This indicates that tasks requiring more complex reasoning benefit from a larger rank, which provides greater expressive capacity. The fact that an optimal rank exists and varies by task highlights that our low-rank decomposition is not merely a compression technique but acts as a crucial, task-adaptive regularizer.
Hallucination analysis on image captioning.

### 26 Mar 2026 v3

@@ Abstract
In-Context Learning (ICL) is a significant paradigm for Large Multimodal Models (LMMs), using a few in-context demonstrations (ICDs) for new task adaptation. However, its performance is sensitive to demonstration configurations and computationally expensive. Mathematically, the influence of these demonstrations can be decomposed into a dynamic mixture of the standard attention output and the context values. Current approximation methods simplify this process by learning a “shift vector”. Inspired by the exact decomposition, we introduce High-Fidelity In-Context Learning (HiFICL) to more faithfully model the ICL mechanism. HiFICL consists of three key components: 1) a set of “virtual key-value pairs” to act as a learnable context, 2) a low-rank factorization for stable and regularized training, and 3) a simple end-to-end training objective. From another perspective, this mechanism constitutes a form of context-aware Parameter-Efficient Fine-Tuning (PEFT). Extensive experiments show that HiFICL consistently outperforms existing approximation methods on several multimodal benchmarks. The code is available at https://github.com/bbbandari/HiFICL.
@@ 1 Introduction
Figure 1: Overview of the HiFICL framework. (a) The overall training process involves a frozen LMM backbone and layer-wise trainable virtual key-value pairs, optimized via a final task loss. (b) Inside each attention head, these virtual pairs are used to compute the high-fidelity attention output according to our derived formula (Eq. 1). (c) Conceptually, unlike previous linear shift methods that only learn an additive offset ($y=x+b$), HiFICL introduces a dynamic scaling factor ($\alpha$) to modulate the original attention ($y=ax+b$). This non-linear partitioning provides a more faithful fit to the underlying data manifold.
@@
To mitigate these issues, a dominant research direction has focused on approximating the ICL effect. This paradigm, centered on learning a “shift vector,” aims to distill the knowledge from explicit ICDs into a compact representation that can be efficiently injected into the model [16, 40, 36, 19, 26]. While this approach has proven effective in reducing inference costs, we argue that this paradigm is predicated on a simplified and potentially restrictive linear assumption. These methods treat the “shift” as an external, additive component to be learned, focusing on approximating the emergent outcome of the ICL process. This overlooks a more fundamental question: what is the underlying causal mechanism that generates this effect in the first place?
@@
This reframing of the ICL approximation problem aligns with a growing body of research questioning the adequacy of the linear shift model. Recent studies have begun to theorize that ICL itself functions as a form of “inference-time finetuning,” where the model’s behavior is dynamically and temporarily optimized by the provided context, much like a few steps of gradient descent [9, 23]. This perspective, supported by empirical evidence showing a striking similarity between the effects of ICL and finetuning on model confidence [2], suggests that the ICL mechanism is fundamentally a learning process, not a simple vector addition. While these works provide the theoretical and empirical grounding for this hypothesis, our work provides the first concrete instantiation of it as a training-time PEFT method. HiFICL is designed to distill and solidify this transient, inference-time adaptation into a permanent and efficient set of trainable parameters. Viewed through this lens, its mechanism presents a more logically coherent form of adaptation than conventional methods like LoRA [17], which apply static, input-agnostic updates. By directly simulating the foundational process of ICL through learnable, context-aware modulation, HiFICL offers a dynamic and principled approach to model adaptation.
@@ 2.1 ICL in Large Multimodal Models
In-Context Learning (ICL) enables few-shot adaptation without parameter updates in Large Language Models (LLMs) [4, 11]. This paradigm has naturally extended to Large Vision-Language Models (LVLMs), where early works established the viability of interleaved image-text conditioning [1, 3, 20, 41]. However, multimodal ICL introduces unique challenges compared to the text domain [33], as models often neglect visual contexts [5]. First, the high token cost of visual inputs drastically limits context windows and incurs prohibitive computational overhead [30, 46]. Second, cross-modal synergies amplify performance sensitivity to demonstration selection and ordering [25, 32]. Unlike in LLMs, visually similar retrievals may misalign with task semantics in LMMs [48], complicating optimal configuration searches in complex feature spaces [37, 49]. While recent works attempt sequence optimization [28] or attention modulation [27, 24, 6] to mitigate instability, the fundamental hurdles of computational cost and extreme sensitivity remain, strongly motivating the paradigm of ICL approximation.
@@ 2.2 The Linear Shift Approximation
However, recent advances in mechanistic interpretability reveal a fundamental contradiction with this linear paradigm. Research has shown that ICL is not a simple global shift but is executed by specialized circuits. A key discovery is the role of “Induction Heads” [35, 13], attention heads that perform complex pattern matching and retrieval. These mechanisms are believed to be how models implicitly implement learning algorithms like gradient descent [9, 45]. Furthermore, analyses from the perspective of hidden state geometry demonstrate that ICL is a highly non-linear transformation that dynamically reshapes the representation space to enhance class separability and align with decision boundaries [47]. This difference between the linear assumption of current approximation methods and the non-linear reality of the underlying mechanism suggests that relying solely on the linear shift paradigm may restrict the model’s full potential, motivating the exploration of more faithful approximation models.
@@ 2.3 Low-Rank Adaptation for PEFT
The connection between In-Context Learning (ICL) and finetuning is not merely a conceptual analogy but is rooted in the mechanistic underpinnings of the Transformer architecture. A growing body of theoretical research posits that the self-attention mechanism can implicitly simulate optimization algorithms. These studies show that attention heads can identify relevant context examples and aggregate their label information in a manner that mathematically mirrors one or more steps of gradient descent on a prediction loss [9, 23, 38]. This frames ICL as a form of “inference-time” optimization, where the model effectively fine-tunes its internal representations dynamically for each query.
This raises a natural question: how does this implicit, dynamic optimization relate to explicit finetuning? Parameter-Efficient Fine-Tuning (PEFT) [15], notably Low-Rank Adaptation (LoRA) [17], represents the dominant paradigm for the latter, injecting trainable low-rank matrices into frozen layers [10]. However, a core characteristic of LoRA is that its adaptation is static and input-agnostic. This creates a conceptual gap: while ICL performs a dynamic, per-query optimization, LoRA performs a static, global one. Our work is motivated by bridging this gap, aiming to create a PEFT method that is not only parameter-efficient but also retains the dynamic, context-aware adaptation properties inherent to the ICL mechanism.
@@ 3.1 Mathematical Analysis
where $\text{SA}(q,K,V)$ is the standard self-attention output over the query tokens and $V_{D}$ can be viewed as the redefined offset vector basis. The coefficients $\alpha(q,K,K_{D})$ and $\beta(q,K,K_{D})$ are dynamically computed scalar and vector weights, respectively, defined as:
$\alpha=\frac{Z_{2}}{Z_{1}+Z_{2}},\quad\beta=\frac{1}{Z_{1}+Z_{2}}\exp\left(\frac{qK_{D}^{\top}}{\sqrt{d_{k}}}\right)$ (2)
@@ 3.2 HiFICL: Direct Parameterization
Based on the theoretical insights from Sec. 3.1, we introduce HiFICL. The core idea is to directly parameterize the source of ICL, the unknown demonstration key-value pairs $(K_{D},V_{D})$, rather than approximating its ultimate effect. We achieve this by introducing a set of virtual context slots, which manifest directly as “virtual key-value pairs” $(K_{learn},V_{learn})$ within the attention mechanism. To capture the functional specialization of different attention heads [8, 44], our modifications are applied on a per-head basis. We equip each attention head $h$ with an independent set of virtual pairs, $(K_{learn}^{(h)},V_{learn}^{(h)})$, allowing each head to learn the specialized contextual information it requires for its sub-task.
@@
where $n$ denotes the sequence length of these virtual matrices, $K_{A}^{(h)},V_{A}^{(h)}\in\mathbb{R}^{n\times r}$ and $K_{B}^{(h)},V_{B}^{(h)}\in\mathbb{R}^{r\times d_{h}}$, with the rank $r\ll d_{h}$. This design elegantly provides two critical benefits. First, for structural stability, we initialize the $V_{B}^{(h)}$ matrices to zero. This guarantees that the contextual shift term in Eq. (1) is null at the beginning of training, preventing potential gradient explosion and creating a smooth learning trajectory from the base model’s state. Second, the low-rank factorization of $K_{learn}^{(h)}$ acts as a powerful form of structural regularization. It compels the model to learn a compact and generalizable set of “prototype keys,” effectively creating an information bottleneck that mitigates overfitting.
To maximize the potential of this high-fidelity architecture, we adopt a simple yet powerful teacher-free, end-to-end optimization strategy. We depart from the teacher-student paradigm used in prior work [19], which relies on intermediate alignment losses. Instead, we optimize all trainable parameters, including the injected low-rank virtual matrices, using only the final supervised task loss. For generative tasks, this is typically the cross-entropy loss over the target sequence:
@@ 3.4 HiFICL as a Dynamic PEFT
Beyond ICL approximation, HiFICL introduces a new perspective on Parameter-Efficient Fine-Tuning (PEFT). As theorized, ICL functions as “inference-time finetuning,” dynamically optimizing model behavior via context [23, 2]. Our work provides the first concrete instantiation of this hypothesis as a training-time PEFT method. Contrasting HiFICL with Low-Rank Adaptation (LoRA) [17], a dominant technique that performs a static, input-agnostic adaptation in the weight space (Fig. 3), HiFICL operates in the activation space, performing a dynamic, content-aware adaptation by providing the model with a learnable memory of virtual contextual examples. Its mechanism is thus more interpretable, analogous to teaching a model how to reason with context rather than applying static, input-agnostic modifications to its weights. As our experiments will demonstrate, this principled design, which effectively distills the “inference-time finetuning” mechanism into a permanent set of parameters, not only leads to superior performance but also achieves significantly higher parameter efficiency.
@@ 4.1 Setup
To ensure a fair comparison, our implementation details for all trainable baselines strictly follow those of MimIC [19]. We employ the AdamW optimizer [31] with a learning rate of 5e-3, coupled with a cosine annealing scheduler with a warmup phase over the first 10% of training steps. For HiFICL, we set the sequence length of the virtual matrices to $n=8$ (mimicking an 8-shot setting) across all experiments. We generally adopt a rank of $r=8$ as a robust default (with occasional adjustments to 4 or 16 depending on dataset complexity), leaving exact configurations detailed in the Appendix.
@@ 4.2 Comparison with Existing Methods
Compared methods.
We compare HiFICL against several categories of strong baselines. (1) In-Context Learning (ICL): We evaluate the performance of Zero-shot and 8-shot ICL with randomly selected demonstrations, and additionally include 16-shot and 32-shot settings on the Idefics2 model for a broader ICL comparison. (2) PEFT: We include LoRA [17], applying it with a rank of 16 to the attention matrices ($W_{q},W_{k},W_{v},W_{o}$) to strictly follow the official setting of MimIC [19] for a direct comparison. (3) ICL Approximation: We benchmark against state-of-the-art methods from this paradigm which modify the language model part of the LMM. These methods differ in their insertion points: LIVE [36] introduces learnable vectors after each FFN layer, whereas both the prior state-of-the-art, MimIC [19], and our HiFICL operate directly within the Multi-Head Attention (MHA) block of each layer. For all trainable methods, we ensure a fair comparison by using the same training data and core hyperparameter settings.
Results analysis.
[…]
@@
We further analyze learning efficiency across five data scales (100 to 1000 samples). As illustrated in Fig. 4, HiFICL consistently demonstrates strong data efficiency. While minor performance fluctuations appear in the extreme low-data regime due to typical few-shot instability, HiFICL generally exhibits a steeper learning curve than competing methods. Notably, we observed that the LoRA baseline suffers from performance degradation in certain scenarios; this is primarily attributed to overfitting caused by the high rank setting ($r=16$) inherited from prior works. Despite this, HiFICL surpasses the strong 8-shot ICL performance with only around 300 training samples (e.g., on Idefics2 for COCO), suggesting it provides a more direct and effective learning signal.
@@ 4.3 Ablations and More Analyses
Core component ablations.
We systematically deconstruct HiFICL on the Idefics2 model to validate its key design principles, with results detailed in Tab. 2. The findings confirm the necessity of each component. First, re-implementing HiFICL within a teacher-student framework (+ Teacher) by aligning hidden states similar to MimIC [19] leads to a consistent performance drop (e.g., VQAv2 accuracy falls from 72.08% to 70.09%), supporting our hypothesis that the teacher model acts as a performance ceiling. Second, the dual low-rank decomposition is vital. While removing the factorization from either the key matrix (- LoRA on K) or the value matrix (- LoRA on V) significantly degrades performance, ablating $V$ incurs a more severe drop than ablating $K$ (e.g., 69.31% vs. 70.58% on VQAv2). This aligns perfectly with our theoretical derivation in Sec. 3.1: $K$ primarily computes the dynamic mixture coefficients, whereas $V$ directly constitutes the contextual shift basis ($V_{D}$), making its parameterization more critical. Finally, disabling the self-attention scaling factor ($\alpha=1$), which reverts our method to a simpler linear shift approximation, also causes a substantial performance decline across all benchmarks. This empirically validates that capturing the full, non-linear attention dynamic is a key driver of HiFICL’s superiority.
@@
A primary motivation for ICL approximation is to reduce the high inference cost of long contexts. We evaluate this by comparing the average processing time per sample on VQAv2 using the LLaVA-Interleave-7b model. As shown in Fig. 6, all ICL approximation methods, including HiFICL, offer a dramatic speedup over standard multi-shot ICL. HiFICL’s inference runtime is nearly on par with the Zero-shot baseline and comparable to prior approximation methods. Crucially, it is approximately 2.0x faster than 8-shot ICL and 3.4x faster than 16-shot ICL, confirming that our high-fidelity approach retains the core efficiency benefits of ICL approximation.
@@
Task-adaptive rank.
Our analysis of the low-rank decomposition on Idefics2 reveals that the optimal rank $r$ correlates closely with task complexity. As shown in Fig. 7, the knowledge-intensive OK-VQA benefits from a higher capacity ($r=16$), whereas VQAv2 peaks earlier at $r=8$. Crucially, HiFICL shares rank-setting principles with LoRA. Empirically, $r=8$ emerges as a robust “sweet spot” across most scenarios, offering a reliable heuristic without extensive search. This demonstrates that our low-rank design acts not merely as a compression technique, but as a crucial, task-adaptive regularizer.
Hallucination analysis on image captioning.